# Restructuring Tractable Probabilistic Circuits

### v1

@@ Appendix C MULTIPLICATION WITH NON-STRUCTURED CIRCUITS
Given a contiguous structured PC $\mathscr{A}$ respecting a linear vtree and an arbitrary contiguous PC $\mathscr{B}$, which is not necessarily structured, we show a recursive algorithm that multiplies $\mathscr{A}$ and $\mathscr{B}$ in polynomial time. Specifically, for each possible scope $\mathbf{X}_{a:b}$ that appears in $\mathscr{B}$, we recursively construct circuit representations for the functions $p_{q}(\mathbf{X}_{a:b})\cdot p_{\mathscr{A}}(\mathbf{X}_{a:b}\>|\>Z_{a}\!=\!i,Z_{b}\!=\!j)$ for $\oplus$ nodes $q\!\in\!\mathscr{B}$ with scope $\mathbf{X}_{a:b}$ and $i,j$ hidden states of $\mathscr{A}$. In particular, from $p_{\mathscr{A}}(\mathbf{X}_{a:b}\>|\>Z_{a}\!=\!i,Z_{b}\!=\!j)$ we drop $Z_{a}=i$ if $a=1$ and drop $Z_{b+1}=j$ if $b=n$, thus $p_{\mathscr{A}}(\mathbf{X}_{a:b}\>|\>Z_{a}\!=\!i,Z_{b}\!=\!j)$ corresponds to $p_{\mathscr{A}}(\mathbf{X}_{a:b}\>|\>\mathbf{C}_{a:b})$ as defined in Case 1. of Section 4.
The recurrence relation is similar to that of Algorithm 1. In the following derivation, we use the notations: (1) denote the children of $\oplus$ node $q$ by $c\!\in\!\textsc{Ch}(q)$; (2) denote the children of $\otimes$ node $c$ by $c_{1}$ and $c_{2}$; (3) denote the weight of the edge connecting $q$ and $c$ by $w_{qc}$; (4) for each $\otimes$ node $c\!\in\!\textsc{Ch}(q)$, $c$ splits $\mathbf{X}_{q}\!=\!\{X_{a},\dots,X_{b}\}$ into two contiguous segments, and denote them by $\mathbf{X}_{c_{1}}\!=\!\{X_{a},\dots,X_{m_{c}}\}$ and $\mathbf{X}_{c_{2}}=\{X_{m_{c}+1},\dots,X_{b}\}$ for some $a\!\leq\!m_{c}\!\leq\!b$.
$\boxed{p_{q}(\mathbf{X}_{a:b})\cdot p_{\mathscr{A}}(\mathbf{X}_{a:{b+1}}\>|\>Z_{a}\!=\!i,Z_{b+1}\!=\!j)}$
$=\sum_{c\in\textsc{Ch}(q)}p_{c}(\mathbf{X}_{a:b})\cdot w_{qc}\cdot p_{\mathscr{A}}(\mathbf{X}_{a:{b+1}}\>|\>Z_{a}\!=\!i,Z_{b+1}\!=\!j)$
$=\sum_{c\in\textsc{Ch}(q)}p_{c_{1}}(\mathbf{X}_{a:m_{c}})\cdot p_{c_{2}}(\mathbf{X}_{m_{c}+1:{b+1}})\cdot w_{qc}\cdot p_{\mathscr{A}}(\mathbf{X}_{a:{b+1}}\>|\>Z_{a}\!=\!i,Z_{b+1}\!=\!j)$
$=\sum_{c\in\textsc{Ch}(q)}p_{c_{1}}(\mathbf{X}_{a:m_{c}})\cdot p_{c_{2}}(\mathbf{X}_{m_{c}+1:b+1})\cdot w_{qc}\cdot\sum_{k}p_{\mathscr{A}}(\mathbf{X}_{a:{b+1}},{Z_{m_{c}}}\!=\!k\>|\>Z_{a}\!=\!i,Z_{b+1}\!=\!j)$
@@
$\quad\cdot p_{c_{1}}(\mathbf{X}_{a:m_{c}})\cdot p_{\mathscr{A}}(\mathbf{X}_{a:m_{c}}\>|\>Z_{m_{c}+1}\!=\!k,Z_{a}\!=\!i,Z_{b+1}\!=\!j)\cdot p_{c_{2}}(\mathbf{X}_{m_{c}+1:b})\cdot p_{\mathscr{A}}(\mathbf{X}_{a:b}\>|\>Z_{m_{c}+1}\!=\!k,Z_{a}\!=\!i,Z_{b+1}\!=\!j)$
@@
$\quad\quad\cdot\boxed{p_{c_{1}}(\mathbf{X}_{a:m_{c}})\cdot p_{\mathscr{A}}(\mathbf{X}_{a:m_{c}}\>|\>Z_{a}\!=\!i,Z_{m_{c}+1}\!=\!k)}\cdot\boxed{p_{c_{2}}(\mathbf{X}_{m_{c}+1:b})\cdot p_{\mathscr{A}}(\mathbf{X}_{a:b}\>|\>Z_{m_{c}+1}\!=\!k,Z_{b+1}\!=\!j)}$
Now let’s analyze the complexity of the constructed circuit, which we denote by $\mathscr{C}$. $\mathscr{C}$ has $O(mkh^{2})$ $\oplus$ nodes in total, where $m$ is the number of scopes in $\mathscr{B}$, $k:=\max_{\mathbf{S}\text{ a scope in $\mathscr{B}$}}|\{\oplus\in\mathscr{B}\text{ with scope }\mathbf{S}\}|$, and $h$ is the hidden states size of $\mathscr{A}$. Suppose that each $\oplus$ node in $\mathscr{B}$ has at most $r$ children, then each $\oplus$ node in $\mathscr{C}$ has at most $O(rh)$ children. Hence the size of $\mathscr{C}$ is bounded by $O(mkh^{2}\cdot rh)=O(mkr\cdot h^{3})$. Note that $O(mkr)$ corresponds to $O(|\mathscr{B}|)$ and $O(h^{3})$ is upper-bounded by $O(h^{4})$, which is $O(|\mathscr{A}|^{2})$; hence the size of $\mathscr{C}$ is bounded by $O(|\mathscr{A}|^{2}|\mathscr{B}|)$, which is the same complexity as stated in Theorem 4.3. Hence, we can remove the assumption that $\mathscr{B}$ has to be structured from Theorem 4.3.
Theorem C.1.
By a similar recursive construction, we can also remove the assumption that $\mathscr{B}$ has to be structured from Theorem 4.5:

### v2

@@ Appendix C MULTIPLICATION WITH NON-STRUCTURED CIRCUITS
Given a contiguous structured PC $\mathscr{A}$ respecting a linear vtree and an arbitrary contiguous PC $\mathscr{B}$, which is not necessarily structured, we show a recursive algorithm that multiplies $\mathscr{A}$ and $\mathscr{B}$ in polynomial time. Specifically, for each possible scope $\bm{X}_{a:b}$ that appears in $\mathscr{B}$, we recursively construct circuit representations for the functions $p_{q}(\bm{X}_{a:b})\cdot p_{\mathscr{A}}(\bm{X}_{a:b}\>|\>Z_{a}\!=\!i,Z_{b}\!=\!j)$ for $\oplus$ nodes $q\!\in\!\mathscr{B}$ with scope $\bm{X}_{a:b}$ and $i,j$ hidden states of $\mathscr{A}$. In particular, from $p_{\mathscr{A}}(\bm{X}_{a:b}\>|\>Z_{a}\!=\!i,Z_{b}\!=\!j)$ we drop $Z_{a}=i$ if $a=1$ and drop $Z_{b+1}=j$ if $b=n$, thus $p_{\mathscr{A}}(\bm{X}_{a:b}\>|\>Z_{a}\!=\!i,Z_{b}\!=\!j)$ corresponds to $p_{\mathscr{A}}(\bm{X}_{a:b}\>|\>\bm{C}_{a:b})$ as defined in Case 1. of Section 4.
The recurrence relation is similar to that of Algorithm 1. In the following derivation, we use the notations: (1) denote the children of $\oplus$ node $q$ by $c\!\in\!\textsc{Ch}(q)$; (2) denote the children of $\otimes$ node $c$ by $c_{1}$ and $c_{2}$; (3) denote the weight of the edge connecting $q$ and $c$ by $w_{qc}$; (4) for each $\otimes$ node $c\!\in\!\textsc{Ch}(q)$, $c$ splits $\bm{X}_{q}\!=\!\{X_{a},\dots,X_{b}\}$ into two contiguous segments, and denote them by $\bm{X}_{c_{1}}\!=\!\{X_{a},\dots,X_{m_{c}}\}$ and $\bm{X}_{c_{2}}=\{X_{m_{c}+1},\dots,X_{b}\}$ for some $a\!\leq\!m_{c}\!\leq\!b$.
$\boxed{p_{q}(\bm{X}_{a:b})\cdot p_{\mathscr{A}}(\bm{X}_{a:{b+1}}\>|\>Z_{a}\!=\!i,Z_{b+1}\!=\!j)}$
$=\sum_{c\in\textsc{Ch}(q)}p_{c}(\bm{X}_{a:b})\cdot w_{qc}\cdot p_{\mathscr{A}}(\bm{X}_{a:{b+1}}\>|\>Z_{a}\!=\!i,Z_{b+1}\!=\!j)$
$=\sum_{c\in\textsc{Ch}(q)}p_{c_{1}}(\bm{X}_{a:m_{c}})\cdot p_{c_{2}}(\bm{X}_{m_{c}+1:{b+1}})\cdot w_{qc}\cdot p_{\mathscr{A}}(\bm{X}_{a:{b+1}}\>|\>Z_{a}\!=\!i,Z_{b+1}\!=\!j)$
$=\sum_{c\in\textsc{Ch}(q)}p_{c_{1}}(\bm{X}_{a:m_{c}})\cdot p_{c_{2}}(\bm{X}_{m_{c}+1:b+1})\cdot w_{qc}\cdot\sum_{k}p_{\mathscr{A}}(\bm{X}_{a:{b+1}},{Z_{m_{c}}}\!=\!k\>|\>Z_{a}\!=\!i,Z_{b+1}\!=\!j)$
@@
$\quad\cdot p_{c_{1}}(\bm{X}_{a:m_{c}})\cdot p_{\mathscr{A}}(\bm{X}_{a:m_{c}}\>|\>Z_{m_{c}+1}\!=\!k,Z_{a}\!=\!i,Z_{b+1}\!=\!j)\cdot p_{c_{2}}(\bm{X}_{m_{c}+1:b})\cdot p_{\mathscr{A}}(\bm{X}_{a:b}\>|\>Z_{m_{c}+1}\!=\!k,Z_{a}\!=\!i,Z_{b+1}\!=\!j)$
@@
$\quad\quad\cdot\boxed{p_{c_{1}}(\bm{X}_{a:m_{c}})\cdot p_{\mathscr{A}}(\bm{X}_{a:m_{c}}\>|\>Z_{a}\!=\!i,Z_{m_{c}+1}\!=\!k)}\cdot\boxed{p_{c_{2}}(\bm{X}_{m_{c}+1:b})\cdot p_{\mathscr{A}}(\bm{X}_{a:b}\>|\>Z_{m_{c}+1}\!=\!k,Z_{b+1}\!=\!j)}$
Now let’s analyze the complexity of the constructed circuit, which we denote by $\mathscr{C}$. $\mathscr{C}$ has $O(mkh^{2})$ $\oplus$ nodes in total, where $m$ is the number of scopes in $\mathscr{B}$, $k:=\max_{\bm{S}\text{ a scope in $\mathscr{B}$}}|\{\oplus\in\mathscr{B}\text{ with scope }\bm{S}\}|$, and $h$ is the hidden states size of $\mathscr{A}$. Suppose that each $\oplus$ node in $\mathscr{B}$ has at most $r$ children, then each $\oplus$ node in $\mathscr{C}$ has at most $O(rh)$ children. Hence the size of $\mathscr{C}$ is bounded by $O(mkh^{2}\cdot rh)=O(mkr\cdot h^{3})$. Note that $O(mkr)$ corresponds to $O(|\mathscr{B}|)$ and $O(h^{3})$ is upper-bounded by $O(h^{4})$, which is $O(|\mathscr{A}|^{2})$; hence the size of $\mathscr{C}$ is bounded by $O(|\mathscr{A}|^{2}|\mathscr{B}|)$, which is the same complexity as stated in Theorem 4.4. Hence, we can remove the assumption that $\mathscr{B}$ has to be structured from Theorem 4.4. See 4.9 By a similar recursive construction, we can also remove the assumption that $\mathscr{B}$ has to be structured from Theorem 4.6:
@@
Appendix D COMPLETE EXAMPLE OF RESTRUCTURING
In the section, we show a simple example of our restructuring algorithm in its entirety. In particular, we will consider the restructuring of a 4-variable circuit following a linear vtree, to a contiguous log-depth vtree.
In Figure 5 we show the procedure of converting the linear vtree into a Bayesian network representation, the target contiguous vtree $W$, and the labeling constructed by Algorithm 2. In particular, this is a valid cover according to Definition 3.7 as $\{Z_{2}\}$ is a cover for both $\{X_{1},X_{2}\}$ and $\{X_{3},X_{4}\}$, and the second and third conditions in the definition are satisfied as e.g. $\{Z_{2}\}$ blocks paths between $\{X_{1},X_{2}\}$ and $\{Z_{2}\}$ (trivially).
We further show an explicit example of the restructuring of a PC respecting the right-linear vtree in Figure 6. In the original PC, we have each $Z_{i}\in\{0,1\}$, corresponding to one of the product nodes in each scope. In the restructuring, we follow the generic construction in Figure 2 for each triple of vtree nodes $(w,l,r)$ given the labeling in Figure 5(c), giving rise to the restructured PC. For example, note that the root probabilities $(0.66,0.34)$ correspond to the marginal probability $p(Z_{2})$ in the original PC. One subtlety is that the leaf nodes do not all directly correspond to the leaf nodes in the original PC; for example $p(X_{1}|Z_{2})=\sum_{Z_{1}}p(Z_{1}|Z_{2})p(X_{1}|Z_{1})$ where $p(X_{1}|Z_{1})$ are the leaf nodes with scope $\{X_{1}\}$ in the original circuit; but this is not problematic as we can simply mix the leaf nodes (either explicitly with a sum node, or by constructing new leaf nodes).
(a) Bayesian Network for linear vtree $V_{v\mapsto Z_{v}}$
(b) Target vtree $W$
(c) Labelled target vtree $W$
Figure 5: Example of restructuring labeling.
(a) Original PC
(b) Restructured PC
Figure 6: Example of original and restructured PC for the structures in Figure 5.